OpenStudy (anonymous):

X/X-3 + X/2 =6X/2X-6

OpenStudy (anonymous):

notice that 2x-6 = 2(x-3)

OpenStudy (anonymous):

so 6x/2x-6 = 2(3x)/2(x-3)=3x/x-3

OpenStudy (anonymous):

THANKS ^^

OpenStudy (anonymous):

i hope you know what to do after that

OpenStudy (anonymous):

YEAH THANKS ^.^